### v1

# Coherent risk measures and uniform integrability

Abstract
We establish a profound connection between coherent risk measures, a prominent object in quantitative finance, and uniform integrability, a fundamental concept in probability theory. Instead of working with absolute values of random variables, which is convenient in studying integrability, we work directly with random loses and gains, which have clear financial interpretation. We introduce a technical tool called the folding score of distortion risk measures. The analysis of the folding score allows us to convert some conditions on absolute values to those on gains and losses. As our main results, we obtain three sets of equivalent conditions for uniform integrability. In particular, a set is uniformly integrable if and only if one can find a coherent distortion risk measure that is bounded on the set, but not finite on $L^{1}$.
Keywords: Expected Shortfall, distortion risk measures, folding scores, law invariance, law of large numbers
1 Introduction
Coherent risk measures, introduced by Artzner et al. (1999), have been a cornerstone of quantitative finance. The development on risk measures has shown to be prolific in both academic research and financial regulation; we refer to Föllmer and Schied (2016) and McNeil et al. (2015) for general treatments on risk measures and their applications. In particular, the Value-at-Risk (VaR) and the Expected Shortfall (ES, also known as CVaR, TVaR and AVaR) are the two most important risk measures in the financial industry; see the regulatory document of BCBS (2019) in the banking sector.
Our main goal is to connect coherent risk measures to uniform integrability, an old concept in probability theory, useful in many domains of analysis. As an example of financial relevance, for a sequence of random variables that converges in distribution, e.g., estimated financial models from data that converge to the true model, uniform integrability guarantees convergence of many risk measures, including ES. Continuity of a risk measure with respect to distributional convergence is associated with robustness of the risk measure by Cont et al. (2010). See Krätschmer et al. (2014) and Embrechts et al. (2022) for recent developments on the robustness of risk measures. Uniform integrability is a useful condition in many financial models, especially in the context of martingales; see e.g., Källblad et al. (2018) and Bayraktar et al. (2020).
We will study conditions on values of a coherent risk measure applied to random variables in a set $\mathcal{S}$ that characterize uniform integrability of $\mathcal{S}$. There is a subtle difference for conditions typically used in probability theory and those in the literature of risk measures. Due to the definition of uniform integrability, it is conventional to consider conditions on $|X|$ for $X\in\mathcal{S}$. Suppose that the random variable $X$ represents a financial position (an asset), with its positive realized values representing losses and negative ones representing gains, a convention used by McNeil et al. (2015) (a sign change from Föllmer and Schied (2016)). For a risk measure $\rho$, the value $\rho(X)$ has a concrete interpretation as regulatory capital requirement for a long position in the asset. Similarly, $\rho(-X)$ also has a concrete meaning, as the regulatory capital requirement for a short position in the asset. In an insurance setting, $\rho(X)$ would represent the price of the random loss $X$. On the other hand, $\rho(|X|)$, which evaluates the risk of the absolute value of $X$, does not have a reasonable financial meaning. Recall that monotonicity of $\rho$ has the natural interpretation of “less loss is better", but $X\mapsto\rho(|X|)$ is not monotone and loses such meaning. The quantity $\rho(|X|)$ seems to be only relevant for technical analysis but it does not appear in any economic problems.
For the above reason, we would hope to formulate conditions using $\rho(X)$ and $\rho(-X)$, which are financially interpretable, instead of using $\rho(|X|)$. Moreover, conditions on $\rho(X)$ and $\rho(-X)$ are easier to check for common financial models of $X$ than $|X|$; see McNeil et al. (2015, Chapter 2) for some existing formulas. To formulate conditions on $\rho(X)$ and $\rho(-X)$, we introduce a technical tool called the folding score of distortion risk measures in Section 3, which quantifies the supremum of $\rho(|Y|)/\max\{\rho(Y),\rho(-Y)\}$ over $Y\in L^{1}$. A useful result on the folding score (Theorem 1) allows us to conveniently convert conditions on $\rho(|X|)$ to those on $\rho(X)$ and $\rho(-X)$. In particular, we show that for the class of coherent distortion risk measures, boundedness of $\rho(|X|)$ for $X\in\mathcal{S}$ is equivalent to boundedness of $\rho(X)$ and $\rho(-X)$, unless $\rho$ is the expectation.
With the help from the folding score, we obtain in Section 4 three different sets of equivalent conditions for uniform integrability, using ES, coherent distortion risk measures, or law-invariant coherent risk measures. As a particularly convenient result, in Theorem 3 we find that $\mathcal{S}$ is uniformly integrable if and only if one can find a coherent distortion risk measure that is not finite on $L^{1}$ and bounded on $\mathcal{S}$ and $-\mathcal{S}$. This result closely resembles the classic characterization of uniform integrability via the de la Vallée Poussin criterion; see e.g., Meyer (1966, Theorem T22). As intermediate results, we also obtain that a law-invariant coherent risk measure is finite on $L^{1}$ if and only if it is controlled by a constant times the expectation for positive random variables, and for coherent distortion risk measures this is equivalent to a bounded slope of its distortion function. These intermediate results are closely related to the recent results of Fröhlich and Williamson (2024) in the context of rearrangement-invariant Banach norms.
Two consequences of our main results on a law of large numbers and the convergence of ES values are presented in Section 5, and they direct follow from uniform integrability. Section 6 concludes the paper.
2 Notation and preliminaries
Let $\mathcal{X}$ be the set of all random variables in an atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$. Let $L^{\infty}$ be the set of essentially bounded random variables in $\mathcal{X}$ and $L^{1}$ be the set of integrable ones in $\mathcal{X}$. Further, $L_{+}^{1}$ (resp. $L^{\infty}_{+}$) is the set of nonnegative random variables in $L^{1}$ (resp. $L^{\infty}$). Almost surely equal random variables are treated as identical. We write $X\buildrel\mathrm{d}\over{=}Y$ if $X$ and $Y$ have the same distribution. We identify constant random variables with elements in $\mathbb{R}$. Denote by $x\vee y=\max\{x,y\}$ and $x\wedge y=\min\{x,y\}$ for real values $x$ and $y$.
A risk measure is a functional $\rho:\mathcal{L}\to(-\infty,\infty]$ that is finite on $\mathbb{R}$, where the domain $\mathcal{L}$ is a convex cone of random variables in $\mathcal{X}$ containing $\mathbb{R}$; examples of $\mathcal{L}$ are $L^{0}$, $L^{1}$ or $L^{\infty}$. We take the interpretation that a random variable in $\mathcal{L}$ represents loss/profit of a financial position. A coherent risk measure is a risk measure $\rho$ satisfying the following four properties (called axioms in the literature).
  • Monotonicity: $\rho(X)\leqslant\rho(Y)$ for all $X,Y\in\mathcal{L}$ with $X\leqslant Y$.
  • Cash invariance: $\rho(X+c)=\rho(X)+c$ for all $X\in\mathcal{L}$ and $c\in\mathbb{R}$.
  • Convexity: $\rho(\lambda X+(1-\lambda)Y)\leqslant\lambda\rho(X)+(1-\lambda)\rho(Y)$ for all $X,Y\in\mathcal{L}$ and $\lambda\in[0,1]$.
  • Positive homogeneity: $\rho(\lambda X)=\lambda\rho(X)$ for all $X\in\mathcal{L}$ and $\lambda>0$.
Moreover, a convex risk measure is a risk measure that satisfies the first three properties in the above list (Föllmer and Schied (2002) and Frittelli and Rosazza Gianin (2002)), and a monetary risk measure is a risk measure that satisfies the first two properties in the above list. We refer to Föllmer and Schied (2016) for interpretations of these properties, which are nowadays standard in the field. Two further properties below are also standard and useful in this paper.
  • Law invariance: $\rho(X)=\rho(Y)$ for all $X,Y\in\mathcal{L}$ with $X\buildrel\mathrm{d}\over{=}Y$.
  • Lower semicontinuity (for $L^{1}$): $\liminf_{n\to\infty}\rho(X_{n})\geqslant\rho(X)$ if $X_{n}\to X$ in $L^{1}$ as $n\to\infty$.
Two popular classes of law-invariant monetary risk measures in financial practice are VaR and ES. VaR at level $p\in(0,1)$ is defined as
$\mathrm{VaR}_{p}(X)=\inf\{x\in\mathbb{R}:\mathbb{P}(X\leqslant x)\geqslant p\},~{}~{}~{}X\in\mathcal{X},$
and ES at level $p\in(0,1)$ as
$\mathrm{ES}_{p}(X)=\frac{1}{1-p}\int_{p}^{1}\mathrm{VaR}_{q}(X)\mathrm{d}q,~{}~{}~{}X\in\mathcal{X}.$
Note that $\mathrm{ES}_{p}(X)$ is finite if and only if $\mathbb{E}[X_{+}]<\infty$. Either VaR or ES can be characterized by a few axioms; see Kou and Peng (2016) and Wang and Zitikis (2021).
Both VaR and ES belong to the more general class of distortion risk measures, which plays an important role in this paper. Let $\mathcal{D}$ be the set of functions $h:[0,1]\to[0,1]$ that are increasing (in the non-strict sense) with $h(0)=0$ and $h(1)=1$. For $h\in\mathcal{D}$, define a distortion risk measure $\rho_{h}$ by
$\rho_{h}(X)=\int_{0}^{\infty}h(\mathbb{P}(X>x))\mathrm{d}x+\int_{-\infty}^{0}(h(\mathbb{P}(X>x))-1)\mathrm{d}x,~{}~{}~{}~{}X\in\mathcal{L},$
where $\mathcal{L}$ is chosen such that $\infty$-$\infty$ - ∞ does not appear above, so that $\rho_{h}(X)$ is well defined for $X\in\mathcal{L}$. The function $h$ is called the distortion functions of $\rho_{h}$. The following properties of distortion risk measures are well known (see Wang et al. (2020) for a summary) and will be used frequently in the paper.
  1. (a)
    The risk measure $\rho_{h}$ is always well defined and finite on $L^{\infty}$.
  2. (b)
[…]
  3. (c)
    If $h$ is left-continuous, then we can write
[…]
  4. (d)
    If $h$ is concave, then $\rho_{h}$ is well defined on $L^{1}$ and lower semicontinuous.
  5. (e)
    The risk measure $\rho_{h}$ is coherent if and only if $h$ is concave.
  6. (f)
    If $h$ is the identity on $[0,1]$, then $\rho_{h}=\mathbb{E}$.
  7. (g)
    If $h(t)=(t/(1-p))\wedge 1$ for $p\in(0,1)$, then $\rho_{h}=\mathrm{ES}_{p}$, which is a coherent risk measure.
For all coherent distortion risk measures in this paper, their domain will be chosen as $L^{1}$.
3 Folding score
3.1 Folding score of distortion risk measures
In what follows, we set $\infty/\infty=1$, $0/0=1$, and $1/0=\infty$, and let $\rho$ be a risk measure on $\mathcal{L}$. A new quantity $s_{\rho}$ will be useful in our later analysis, which is defined by
$s_{\rho}=\sup_{X\in\mathcal{L}}\frac{\rho(|X|)}{|\rho(X)|\vee|\rho(-X)|}=\sup_{X\in\mathcal{L}}s_{\rho}(X),\mbox{~{}~{}~{}where }s_{\rho}(X)=\frac{\rho(|X|)}{|\rho(X)|\vee|\rho(-X)|}.$
The quantity $s_{\rho}$ will be called the folding score of $\rho$. To explain the name, recall that the distribution of $|X|$ is the folded distribution of $X$. Intuitively, $s_{\rho}$ measures how large $\rho(|X|)$, the risk measure $\rho$ applied to the “folded" random variable $X$, can be relative to $|\rho(X)|$ and $|\rho(-X)|$. Clearly, $s_{\rho}\geqslant 1$ for common risk measures, e.g., those satisfying $\rho(x)=x$ for $x\in\mathbb{R}$. We are particularly interested in whether $s_{\rho}$ is finite. If $\rho=\mathbb{E}^{Q}$ for a probability $Q$ (absolutely continuous with respect to $\mathbb{P}$), then $s_{\rho}$ is infinite, because the denominator can be 00 while the numerator is positive; for instance, this happens for any non-degenerate random variable symmetric about 00. Therefore, for a linear functional $\rho=\mathbb{E}^{Q}$, it is $s_{\rho}=\infty$.
If $\rho$ is a convex risk measure with $\rho(0)=0$, then $\rho(X)+\rho(-X)\geqslant 2\rho(0)=0$ for all $X\in\mathcal{L}$. Therefore, we can remove the absolute values in the denominator in the definition of $s_{\rho}$ and $s_{\rho}(X)$, that is,
[…]
We will conveniently work with this formulation since most of our results are obtained for coherent risk measures.
Our main result in this section is an upper bound on $s_{\rho}$ for a coherent distortion risk measure $\rho$. In particular, we would like to understand whether $\rho=\mathbb{E}$ is the only case where $s_{\rho}=\infty$ among distortion risk measures. This turns out to be true.
Theorem 1.
Suppose that $\rho$ is a coherent distortion risk measure on $L^{1}$ with distortion function $h\in\mathcal{D}$. If $h$ is not the identity, then
$1\leqslant s_{\rho}\leqslant\frac{h(1/2)+1/2}{h(1/2)-1/2}<\infty.$ (1)
As a consequence, $s_{\rho}=\infty$ if and only if $\rho$ is the mean.
Before proving the theorem, we first present a simple algebraic lemma.
Lemma 1.
For $a,b\in[0,1]$, we have
$\max_{x,y>0}\frac{x+y}{(x-ay)\vee(y-bx)}=\frac{2+a+b}{1-ab}.$
Proof.
With the substitution $z=y/x$,
$\max_{x,y>0}\frac{x+y}{(x-ay)\vee(y-bx)}=\max_{z>0}\frac{1+z}{(1-az)\vee(z-b)}.$
It suffices to maximize $f(z)$ defined by $f(z)=({1+z})/({(1-az)\vee(z-b)})$ for $z>0$. If $ab=1$ (i.e., $a=b=1$), then by taking $z=1$ we have $f(z)=\infty$. Next, suppose $ab<1$. Let $z_{0}=(1+b)/(1+a)$. Note that $ab<1$ implies $z_{0}>b$, and we have
$f(z_{0})=\frac{2+a+b}{1-ab}.$
If $z\leqslant z_{0}$ then $f(z)=(1+z)/(1-az)$, which is increasing in $z$, yielding $f(z)\leqslant f(z_{0})$. If $z\geqslant z_{0}$, then $f(z)=(1+z)/(z-b)=(1-b)/(z-b)+1$, which is decreasing in $z$, yielding $f(z)\leqslant f(z_{0})$. Therefore, $z=z_{0}$ maximizes $f(z)$. ∎
Proof of Theorem 1.
Note that if $h$ is concave and not the identity, then $h(q)>q$ for all $q\in(0,1)$, and hence the third inequality in (1) is automatic. The first inequality $1\leqslant s_{\rho}$ is also automatic. Below, we show the remaining inequality in (1), that is,
$\frac{\rho_{h}(|X|)}{\rho_{h}(X)\vee\rho_{h}(-X)}\leqslant\frac{h(1/2)+1/2}{h(1/2)-1/2}\mbox{~{}~{}~{}for all $X\in L^{1}$.}$
Take $X\in L^{1}$. If $X\geqslant 0$ or $X\leqslant 0$, then ${\rho_{h}(|X|)}={\rho_{h}(X)\vee\rho_{h}(-X)},$ and the desired inequality holds. For a general $X$, define the following quantities,
$A$ $=\int_{0}^{\infty}h(\mathbb{P}(X>x))\mathrm{d}x,$
$B$ $=\int_{0}^{\infty}h(\mathbb{P}(-X>x))\mathrm{d}x=\int_{-\infty}^{0}h(\mathbb{P}(X<x))\mathrm{d}x,$
$C$ $=\int_{-\infty}^{0}(1-h(\mathbb{P}(X>x)))\mathrm{d}x,$
$D$ $=\int_{-\infty}^{0}(1-h(\mathbb{P}(-X>x)))\mathrm{d}x=\int_{0}^{\infty}(1-h(\mathbb{P}(X<x)))\mathrm{d}x.$
Note that $A,B,C,D$ are all nonnegative, and $C$ and $D$ are bounded from above. Clearly,
$\rho_{h}(X)=A-C~{}~{}\mbox{and}~{}~{}\rho_{h}(-X)=B-D.$
Moreover, concavity of $h$ implies $h(s+t)=h(s+t)+h(0)\leqslant h(s)+h(t)$ for $s,t\in[0,1]$ and $s+t\leqslant 1$. Since $\mathbb{P}(|X|>x)=\mathbb{P}(X>x)+\mathbb{P}(X<-x)$ for $x\geqslant 0$, we have
$\rho_{h}(|X|)$ $=\int_{0}^{\infty}h(\mathbb{P}(X>x)+\mathbb{P}(X<-x))\mathrm{d}x$
$\leqslant\int_{0}^{\infty}h(\mathbb{P}(X>x))\mathrm{d}x+\int_{0}^{\infty}h(\mathbb{P}(X<-x))\mathrm{d}x=A+B.$
If one of $A$ and $B$ is infinite, then $\rho_{h}(|X|)\geqslant A\vee B=\infty=\rho_{h}(X)\vee\rho_{h}(-X)$, and the desired inequality (1) holds. Below we assume that $A$ and $B$ are finite.
Let $g\in\mathcal{D}$ be defined by $g(t)=1-h(1-t)$. Clearly, $g$ is convex. Noting that $t\mapsto g(t)/t$ is increasing and $t\mapsto h(t)/t$ is decreasing, we have that $g/h$ is increasing on $(0,1]$. Therefore, using the fact that $\mathbb{P}(X=x)>0$ for at most countably many $x\in\mathbb{R}$, we have
$\frac{D}{A}$ $=\frac{\int_{0}^{\infty}(1-h(\mathbb{P}(X<x)))\mathrm{d}x}{\int_{0}^{\infty}h(\mathbb{P}(X>x))\mathrm{d}x}$
$=\frac{\int_{0}^{\infty}g(\mathbb{P}(X>x))\mathrm{d}x}{\int_{0}^{\infty}h(\mathbb{P}(X>x))\mathrm{d}x}\leqslant\sup_{x\geqslant 0}\frac{g(\mathbb{P}(X>x))}{h(\mathbb{P}(X>x))}.$
Let $z=\mathbb{P}(X<0)$, we have
$\frac{g(\mathbb{P}(X>x))}{h(\mathbb{P}(X>x))}\leqslant\frac{g(\mathbb{P}(X>0))}{h(\mathbb{P}(X>0))}=\frac{g(1-z)}{h(1-z)}=:b,$
and therefore $D/A\leqslant b$. Analogous arguments also yield
$\frac{C}{B}\leqslant\frac{\int_{-\infty}^{0}g(\mathbb{P}(X<x))\mathrm{d}x}{\int_{-\infty}^{0}h(\mathbb{P}(X<x))\mathrm{d}x}\leqslant\frac{g(z)}{h(z)}=:a.$
Note that both $a$ and $b$ take values in $[0,1)$ since $g$ is convex and not the identity. The above inequalities yield
$\frac{\rho_{h}(|X|)}{\rho_{h}(X)\vee\rho_{h}(-X)}\leqslant\frac{A+B}{(A-C)\vee(B-D)}\leqslant\frac{A+B}{(A-aB)\vee(B-bA)}.$
Using Lemma 1, we get
$s_{\rho_{h}}(X)\leqslant\frac{2+a+b}{1-ab}.$
Note that $a$ increases in $z$ and $b$ decreases in $z$. Denote by $c=g(1/2)/h(1/2)$. Since either $z\leqslant 1/2$ or $z>1/2$, we have either $a\leqslant c$ or $b\leqslant c$. This gives $a+b\leqslant 1+c$ and $ab\leqslant c$. Therefore,
$s_{\rho_{h}}(X)\leqslant\frac{3+c}{1-c}=\frac{3-2g(1/2)}{1-2g(1/2)}=\frac{h(1/2)+1/2}{h(1/2)-1/2}.$
The desired inequality (1) follows. ∎
Note that in Theorem 1, concavity of $h$ does not exclude the possibility that $h$ is discontinuous at 00. The following proposition states a simple consequence of $s_{\rho}<\infty$. Our later results will mainly use this proposition.
Proposition 1.
Let $\mathcal{S}$ be a set of random variables and $\rho$ be a convex risk measure satisfying $\rho(0)=0$ and $s_{\rho}<\infty$. The following are equivalent.
  1. (i)
    $\rho(X)$ and $\rho(-X)$ are both bounded for $X\in\mathcal{S}$;
  2. (ii)
    $\rho(|X|)$ is bounded for $X\in\mathcal{S}$.
In particular, this equivalence holds for $\rho=\rho_{h}$, where $h\in\mathcal{D}$ is concave and not the identity.
Proof.
(i)$\Rightarrow$(ii). This follows from the definition of $s_{\rho}$.
(ii)$\Rightarrow$(i). Since the risk measure $\rho$ is monotone, we have $\rho(|X|)\geqslant\rho(X)$ and $\rho(|X|)\geqslant\rho(-X)$; hence, both $\rho(X)$ and $\rho(-X)$ are bounded from above. Convexity of $\rho$ gives $\rho(X)+\rho(-X)\geqslant 2\rho(0)=0$. This guarantees that $\rho(X)$ and $\rho(-X)$ are bounded from below, noting that each of $\rho(X)$ and $\rho(-X)$ is bounded from above.
The last statement for $\rho=\rho_{h}$ follows from Theorem 1. ∎
In Section 4, we will develop a similar pair of conditions involving $|X|$ and ±$\pm X$, both being a characterization of uniform integrability.
Remark 1.
Let $\rho$ be a law-invariant coherent risk measure. If
[…]
then for $Y=X-\rho(X)$, we have $\rho(Y)=0=\rho(-Y)$, but $\rho(|Y|)\geqslant\mathbb{E}[|Y|]>0$. Therefore, $s_{\rho}=\infty$ in this case. Note that (2) is sufficient to force $\rho$ to be the mean; see Castagnoli et al. (2004), Bellini et al. (2021) and Liebrich and Munari (2022) for related results and generalizations. Our Theorem 1 can be seen as a strengthening of this result for distortion risk measures, because $s_{\rho}=\infty$ is a weaker condition than (2).
3.2 Sharpness of the bound in Theorem 1
Let us denote the upper bound in (1) by
$s_{\rho_{h}}\leqslant b_{h}=\frac{h(1/2)+1/2}{h(1/2)-1/2}.$
Note that $b_{h}$ is decreasing in $h(1/2)$, and its smallest value $3$ is attained when $h(1/2)=1$. We provide two examples, first showing that the bound $3$ cannot be improved, and second showing that it is not always sharp for a given $h$.
Example 1 (The upper bound $3$ is sharp for some $h$).
Let $\rho_{h}=\mathrm{ES}_{p}$ for $p\in[1/2,1)$. In this case, $h(1/2)=1$, and the upper bound in (1) is $3$. For $\varepsilon\in(0,1)$, we can construct $X$ such that $s_{\rho_{h}}(X)$ is precisely $3-\varepsilon$, and this shows $s_{\rho_{h}}=3$. Thus, the bound $3$ cannot be improved for $\rho_{h}$. Let $X$ have a two-point distribution given by $\mathbb{P}(X=-1)=1-w$ and $\mathbb{P}(X=2(1-p)/w)=w$, where $w=\varepsilon({1-p})/({4-\varepsilon})\in(0,1-p)$. We have
$\mathrm{ES}_{p}(|X|)=2+(1-p-w)\frac{1}{1-p}=3-\frac{w}{1-p}=3-\frac{\varepsilon}{4-\varepsilon},$
and
$\mathrm{ES}_{p}(X)\vee\mathrm{ES}_{p}(-X)=\left(2-(1-p-w)\frac{1}{1-p}\right)\vee 1=1+\frac{w}{1-p}=1+\frac{\varepsilon}{4-\varepsilon}.$
Therefore,
$\frac{\mathrm{ES}_{p}(|X|)}{\mathrm{ES}_{p}(X)\vee\mathrm{ES}_{p}(-X)}=\frac{3(4-\varepsilon)-{\varepsilon}}{(4-\varepsilon)+\varepsilon}=3-\varepsilon.$
Example 2 (The upper bound in (1) is not sharp for some $h$).
Let $\rho_{h}=\mathrm{ES}_{p}$ for $p\in(0,1/2]$. In this case, $h(1/2)=(2(1-p))^{-1}$, and the upper bound in (1) is $12/p-$ 1. Note that this bound increases to $\infty$ as $p\downarrow 0$. This is intuitive, as $\mathrm{ES}_{0}$ corresponds to the mean, whose folding score is infinity. The bound $12/p-$ 1 is not sharp. Take $p=1/4$, and thus $72/p-1=$ 7. However, for any $z\in(0,1)$, $a+b=g(z)/h(z)+g(1-z)/h(1-z)\leqslant 1$, where $a,b$ are defined in the proof of Theorem 1. Hence, $s_{\rho_{h}}\leqslant(2+a+b)/(1-ab)\leqslant 3/(1-c)=6$.
Although the bound in (1) is not always sharp, it suffices for our study in Section 4, where we only need $s_{\rho}<\infty$ as in Proposition 1.
3.3 Other families of risk measures
Theorem 1 shows that, for the large class of coherent distortion risk measures $\rho$, we have
$s_{\rho}=\sup_{X\in\mathcal{L}}\frac{\rho(|X|)}{|\rho(X)|\vee|\rho(-X)|}<\infty\mbox{ unless $\rho=\mathbb{E}$},$ (3)
One may naturally wonder whether (3) holds for other classes of risk measures, that is, whether the folding score is finite for non-linear risk measures. With a series of counter examples, we answer the question negatively for many classes of interesting risk measures. It suffices to construct examples in $L^{\infty}$, which we will work with in all examples below.
Example 3 (Law-invariant convex risk measures).
Property (3) does not hold for law-invariant convex risk measures in general, illustrated below. The entropic risk measures, indexed by parameter $\beta>0$, are defined by
$\mathrm{ER}_{\beta}(X)=\frac{1}{\beta}\log\mathbb{E}[\exp(\beta X)],~{}~{}~{}~{}~{}X\in L^{1}.$
The entropic risk measures are a popular class of law-invariant convex risk measures (Föllmer and Schied (2016)). We will show that this class does not satisfy (3). Taking $X_{\lambda}=\lambda(2\mathds{1}_{A}-1)$ where $\lambda>0$ and $A\in\mathcal{F}$ with $\mathbb{P}(A)=1/2$, we have $\mathrm{ER}_{\beta}(|X_{\lambda}|)=\lambda$ and
$\mathrm{ER}_{\beta}(X_{\lambda})=\mathrm{ER}_{\beta}(-X_{\lambda})=\frac{1}{\beta}\log\left(\frac{1}{2}\exp(\beta\lambda)+\frac{1}{2}\exp(-\beta\lambda)\right)\geqslant 0.$
Therefore, we have
$\lim_{\lambda\downarrow 0}\frac{\mathrm{ER}_{\beta}(X_{\lambda})}{\mathrm{ER}_{\beta}(|X_{\lambda}|)}=\lim_{\lambda\downarrow 0}\frac{\frac{1}{\beta}\log\left(\frac{1}{2}\exp(\beta\lambda)+\frac{1}{2}\exp(-\beta\lambda)\right)}{\lambda}=\lim_{\lambda\downarrow 0}\frac{\log\left(\frac{1}{2}\exp(\lambda)+\frac{1}{2}\exp(-\lambda)\right)}{\lambda}=0,$
where we used the l’Hospital rule and
$\frac{\mathrm{d}}{\mathrm{d}\lambda}\log\left(\frac{1}{2}\exp(\lambda)+\frac{1}{2}\exp(-\lambda)\right)=\frac{\frac{1}{2}\exp(\lambda)-\frac{1}{2}\exp(-\lambda)}{\frac{1}{2}\exp(\lambda)+\frac{1}{2}\exp(-\lambda)}\to 0\mbox{~{}as $\lambda\downarrow 0$}.$
Therefore, $s_{\rho}=\infty$ for $\rho=\mathrm{ER}_{\beta}$.
Example 4 (Coherent risk measures).
Property (3) does not generalise to coherent, Fatou-continuous risk measures. On $L^{\infty}$, this class of risk measures can be represented by
$\rho=\sup_{Q\in\mathcal{Q}}\mathbb{E}^{Q},$
where $\mathcal{Q}$ is a set of probability measures absolutely continuous with respect to $\mathbb{P}$. Consider $[0,1]$ with the Lebesgue measure $\lambda$. Let $Q_{1}$ and $Q_{2}$ be defined by their Radon-Nikodym derivatives
[…]
Let $X=\mathds{1}_{[0,1/3]}-\mathds{1}_{[2/3,1]},\mathcal{Q}=\{Q_{1},Q_{2}\},$ and $\rho=\sup_{Q\in\mathcal{Q}}\mathbb{E}^{Q}.$ Then $s_{\rho}(X)=1/0=\infty$.
Example 5 (Choquet coherent risk measures).
Property (3) does not hold for the class of Choquet coherent risk measures. A Choquet coherent measure has the form
$\rho_{\nu}(X)=\int_{0}^{\infty}\nu(X>x)\mathrm{d}x+\int_{-\infty}^{0}(\nu(X>x)-1)\mathrm{d}x,~{}~{}~{}X\in L^{\infty},$
where $\nu:\mathcal{F}\to[0,1]$ is increasing and satisfies $\nu(N)=0$ for $N\in\mathcal{F}$ with 00 probability, $\nu(\Omega)=1$, and $\nu(A\cup B)+\nu(A\cap B)\leqslant\nu(A)+\nu(B)$ for all $A,B\in\mathcal{F}$. If $\nu=h\circ\mathbb{P}$ for a concave $h\in\mathcal{D}$, then $\rho_{\nu}=\rho_{h}$. Consider $S=[-1/3,1/3]\times[-1,1]$ with Lebesgue measure $\mu$. Denote $S^{+}=[-1/3,1/3]\times[0,1]$ and $S^{-}=[-1/3,1/3]\times[-1,0)$. Define $\nu(A):=(\mu(A\cap S^{+})\wedge 1/2)+(\mu(A\cap S^{-})\wedge 1/2)$. Take $X=\mathds{1}_{S^{+}}-\mathds{1}_{S^{-}}$. Then $s_{\rho_{\nu}}(X)=1/0=\infty$.
Remark 2.
It remains an open question whether property (3) holds for the class of law-invariant coherent risk measures. For this class, we did not find any example of $\rho\neq\mathbb{E}$ satisfying $s_{\rho}=\infty$, although we could not prove $s_{\rho}<\infty$ for all $\rho\neq\mathbb{E}$. By Kusuoka’s representation (Kusuoka (2001)), any law-invariant coherent risk measure $\rho$ on $L^{\infty}$ can be represented as $\rho=\sup_{h\in\mathcal{H}_{\rho}}\rho_{h}$ where $\mathcal{H}_{\rho}$ is a set of concave distortion functions. If $\mathcal{H}_{\rho}$ is a finite set, then by using Theorem 1 we can obtain (3). For an infinite $\mathcal{H}_{\rho}$, this is not clear.
4 Uniform integrability
4.1 Characterizing uniform integrability using ES
A set $\mathcal{S}$ of random variables is uniformly integrable if
$\sup_{X\in\mathcal{S}}\mathbb{E}[|X|\mathds{1}_{\{|X|>K\}}]\to 0\mbox{~{}as $K\to\infty$}.$
In our case, the random variables in $\mathcal{S}$ are defined on an atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$. Under this setting, uniform integrability of $\mathcal{S}$ has the following characterization (see Meyer (1966, Theorem T19)):
For every ε>0, there exists δ>0 such that $\mbox{For every $\varepsilon>0$,}&\mbox{ there exists $\delta>0$ such that }\\ &\int_{A}|X|\mathrm{d}\mathbb{P}<\varepsilon\mbox{~{}for all $X\in\mathcal{S}$and $A\in\mathcal{F}$ with $\mathbb{P}(A)\leqslant\delta$}.\end{aligned}$ (4)
Moreover, we can safely replace $\mathbb{P}(A)\leqslant\delta$ in (4) by $\mathbb{P}(A)=\delta$.
We have the following characterization theorem of uniform integrability. The equivalence of (i) and (ii) appeared in Wu et al. (2024, Lemma 7.3); here we supply a shorter proof. Statement (ii) below is easier to work with in technical analysis, whereas statement (iii) has clearer meaning for financial applications.
Theorem 2.
Let $\mathcal{S}$ be a set of random variables. The following statements are equivalent.
  1. (i)
    $\mathcal{S}$ is uniformly integrable;
  2. (ii)
    $\sup_{X\in\mathcal{S}}(1-p)\mathrm{ES}_{p}(|X|)$ tends to 00 as $p\uparrow 1$;
  3. (iii)
    both $\sup_{X\in\mathcal{S}}(1-p)\mathrm{ES}_{p}(X)$ and $\sup_{X\in\mathcal{S}}(1-p)\mathrm{ES}_{p}(-X)$ tend to 00 as $p\uparrow 1$.
Proof of Theorem 2.
(i)$\Leftrightarrow$(ii). We will use the following well-known result on the representation of ES (e.g., Eq. (3.1) of Embrechts and Wang (2015)): For any random variable $X$ and $p\in(0,1)$,
$\mathrm{ES}_{p}(X)=\sup\{\mathbb{E}[X|A]:A\in\mathcal{F},~{}\mathbb{P}(A)=1-p\}.$ (5)
Given $\varepsilon>0$, by letting $\delta=1-p$ in (ii), using (5), we have
$\sup_{X\in\mathcal{S}}(1-p)\mathrm{ES}_{p}(|X|)$ $=\sup_{X\in\mathcal{S}}\sup_{A\in\mathcal{F}_{\delta}}\delta\mathbb{E}[X|A]=\sup_{X\in\mathcal{S}}\sup_{A\in\mathcal{F}_{\delta}}\int_{A}|X|\mathrm{d}\mathbb{P},$
where $\mathcal{F}_{\delta}=\{A\in\mathcal{F}:\mathbb{P}(A)=\delta\}$. Since uniform integrability (i) is equivalent to (4), we know that it is also to equivalent to (ii).
(ii)$\Leftrightarrow$(iii). For $p>1/2$ the ratio of $\mathrm{ES}_{p}(|X|)$ to $\mathrm{ES}_{p}(X)\vee\mathrm{ES}_{p}(-X)$ is bounded below by $1$ and above by a constant from Theorem 1 (Example 1 shows this range is contained in $[1,3]$). As $p\uparrow 1$, this uniform bound applies and hence (ii) and (iii) are equivalent. ∎
4.2 Uniform integrability and distortion risk measures
Using the equivalence of (i) and (ii) from Theorem 2, we have the following characterization of uniform integrability in terms of distortion risk measures.
Theorem 3.
Let $\mathcal{D}_{c}$ be the set of concave functions $h\in\mathcal{D}$ with $h(t)/t\to\infty$ as $t\downarrow 0$. For a set $\mathcal{S}$ of random variables, the following are equivalent.
  1. (i)
    $\mathcal{S}$ is uniformly integrable;
  2. (ii)
    there exists $h\in\mathcal{D}_{c}$ such that $\rho_{h}(|X|)$ is bounded for $X\in\mathcal{S}$;
  3. (iii)
    there exists $h\in\mathcal{D}_{c}$ such that $\rho_{h}(X)$ and $\rho_{h}(-X)$ are bounded for $X\in\mathcal{S}$.
Proof.
(ii)$\Rightarrow$(i). We proceed by contrapositive. Suppose that $\mathcal{S}$ is not uniformly integrable. Then there exists $m>0$ and for each $n\in\mathbb{N}$, there exist $p_{n}$ and $X_{n}$ such that $(1-p_{n})\mathrm{ES}_{p_{n}}(|X_{n}|)>m$.
We show that $\rho_{h}(|X|)$ is unbounded for $X\in\mathcal{S}$. That is, given $M>0$, we construct some $X$ such that $\rho_{h}(|X|)>M.$ Choose $p_{N}$ in the sequence $p_{n}$ such that $1-p<M/m$. Let $p=p_{N}$ and $X=X_{N}$. Denote by
$h_{p}(t)=(t/(1-p))\wedge 1,$ (6)
that is, the distortion function corresponding to $\mathrm{ES}_{p}$. We have
$\rho_{h}(|X|)$ $=\int_{0}^{\infty}h(\mathbb{P}(|X|>x))\mathrm{d}x$
$\geqslant\int_{0}^{\infty}h(\mathbb{P}(|X|>x)\wedge(1-p))\mathrm{d}x$
$\geqslant\frac{M}{m}\int_{0}^{\infty}\mathbb{P}(|X|>x)\wedge(1-p)\mathrm{d}x$
$\geqslant\frac{M}{m}\int_{0}^{\infty}(1-p)h_{p}(\mathbb{P}(|X|>x))\mathrm{d}x\geqslant\frac{M}{m}(1-p)\mathrm{ES}_{p}(|X|)\geqslant M.$
Hence, $\{\rho_{h}(|X|):X\in\mathcal{S}\}$ is not bounded.
(i)$\Rightarrow$(ii). Using Theorem 2, (i) implies $\sup_{X\in\mathcal{S}}(1-p)\mathrm{ES}_{p}(|X|)\downarrow 0$ as $p\uparrow 1$. We can choose a sequence $(p_{n})_{n\in\mathbb{N}}\subseteq(0,1)$ such that $1-p_{n}<2^{-n}$ and $\sup_{X\in\mathcal{S}}(1-p_{n})\mathrm{ES}_{p_{n}}(|X|)<2^{-n}$ for each $n$. Define for $t\in[0,1]$, with $h_{p}$ defined in (6),
$g(t)=\sum_{n=1}^{\infty}(1-p_{n})h_{p_{n}}(t)=\sum_{n=1}^{\infty}t\wedge(1-p_{n}).$
It is straightforward to verify that $g$ is finite as $g(t)\leqslant g(1)=\sum_{n=1}^{\infty}(1-p_{n})\leqslant\sum_{n=1}^{\infty}2^{-n}=1$ for $t\in[0,1]$. It is concave, because it is the sum of concave functions. Moreover, $g(0)=0$ and $g(1)>0$. We further define
$h(t)=\frac{g(t)}{g(1)}=\frac{\sum_{n=1}^{\infty}t\wedge(1-p_{n})}{\sum_{n=1}^{\infty}1\wedge(1-p_{n})},$
which is simply the normalized version of $g$ to ensure $h(1)=1$. The properties for $g$ imply that $h$ is a distortion function that is concave. Moreover, we can verify that for any $N\in\mathbb{N}$,
$\lim_{t\downarrow 0}\frac{g(t)}{t}=\lim_{t\downarrow 0}\sum_{n=1}^{\infty}1\wedge\frac{1-p_{n}}{t}\geqslant\lim_{t\downarrow 0}\sum_{n=1}^{N}1\wedge\frac{1-p_{n}}{t}=N.$
Therefore, $\lim_{t\downarrow 0}g(t)/t=\infty.$ Finally,
$\rho_{h}(|X|)$ $=\frac{1}{g(1)}\int_{0}^{\infty}g(\mathbb{P}(|X|>x))\mathrm{d}x$
$=\frac{1}{g(1)}\int_{0}^{\infty}\sum_{n=1}^{\infty}\mathbb{P}(|X|>x)\wedge(1-p_{n})\mathrm{d}x$
$=\frac{1}{g(1)}\sum_{n=1}^{\infty}(1-p_{n})\mathrm{ES}_{p_{n}}(|X|)\leqslant\frac{1}{g(1)}\sum_{n=1}^{\infty}2^{-n}=\frac{1}{g(1)}<\infty.$
Therefore, (ii) holds.
(ii)$\Leftrightarrow$(iii). This follows from Proposition 1. ∎
Theorem 3 closely resembles the classic de la Vallée Poussin criterion on uniform integrability: A set $\mathcal{S}$ is uniformly integrable if and only if there exists an increasing convex function ϕ:++:italic-$\phi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ with $\phi(x)/x\to\infty$ as $x\to\infty$ such that $\sup_{X\in\mathcal{S}}\mathbb{E}[\phi(|X|)]<\infty$; see Meyer (1966, Theorem T22). This highlights a duality between $X\mapsto\rho_{h}(X)$ for a concave $h$ and $X\mapsto\mathbb{E}[\phi(X)]$ for a convex $\phi$. Indeed, $\rho_{h}$ is called the dual utility model of Yaari (1987) in decision theory, and $X\mapsto\mathbb{E}[\phi(X)]$ represents an expected utility model in the same context.
The condition $h\in\mathcal{D}_{c}$ in Theorem 3 allows for $h$ to have a jump at 00, which makes $h(t)/t\to 0$ automatically as $t\downarrow 0$. Such $h$ is not interesting, because boundedness of $\rho_{h}(|X|)$ over $X\in\mathcal{S}$ for such $h$ forces $\mathcal{S}$ to be bounded, an obvious case of uniform integrability.
In part (ii) or (iii) of Theorem 3, $\rho_{h}$ serves as a “testing" risk measure for uniform integrability of $\mathcal{S}$. In the statement of Theorem 3, such a choice of $\rho_{h}$ depends on $\mathcal{S}$. One may wonder whether there is a single $\rho_{h}$ that works for all sets $\mathcal{S}$. To answer this question, we first obtain a result on the finiteness of coherent distortion risk measures, which may be of independent interest. For a risk measure $\rho$ on $L^{1}$, we say that $\rho$ is expectation-dominated if there exists $c>0$ such that $\rho\leqslant c\,\mathbb{E}$ on $L_{+}^{1}$. For instance, $\mathrm{ES}_{p}$ is expectation-dominated for each $p\in(0,1)$, since $\mathrm{ES}_{p}(X)\leqslant(1-p)^{-1}\mathbb{E}[X]$ for $X\in L^{1}_{+}$.
Theorem 4.
Let $\rho$ be a coherent distortion risk measure on $L^{1}$. The following are equivalent.
  1. (i)
    $\rho$ is expectation-dominated;
  2. (ii)
    $\rho$ is finite on $L^{1}$;
  3. (iii)
    the distortion function $h$ of $\rho$ satisfies $\lim_{t\downarrow 0}h(t)/t<\infty$.
Proof.
(i)$\Rightarrow$(ii). Since $\rho\leqslant c\mathbb{E}$ on $L^{1}_{+}$ for some $c>0$, we have $\rho(|X|)<\infty$ for all $X\in L^{1}$. Since $\rho$ is convex, $\rho(-|X|)+\rho(|X|)\geqslant 2\rho(0)=0$, which implies $\rho(-|X|)>-\infty$. Monotonicity of $\rho$ yields $-\infty<\rho(-|X|)\leqslant\rho(X)\leqslant\rho(|X|)<\infty$, showing that $\rho$ is finite on $L^{1}$.
(ii)$\Rightarrow$(iii). Note that $\lim_{t\downarrow 0}h(t)/t$ exists since $h(t)/t$ is decreasing in $t$, due to concavity of $h$. Suppose $h(t)/t=\infty$, and we will show that $\rho$ cannot be finite on $L^{1}$. Let $U$ be uniformly distributed on $[0,1]$. For $t\in(0,1)$, let $B_{t}=\mathds{1}_{\{U<t\}}$ and
$X_{t}=\left(\frac{1}{th(t)}\right)^{1/2}B_{t}.$
Note that $\mathbb{E}[X_{t}]=(t/h(t))^{1/2}\leqslant 1.$ It is straightforward to compute $\rho(X_{t})=(h(t)/t)^{1/2}$ for $t\in(0,1)$, and hence
$\lim_{t\downarrow 0}\rho(X_{t})=\lim_{t\downarrow 0}\left(\frac{h(t)}{t}\right)^{1/2}=\infty.$
Let $t_{n}>0$ be such that $\rho(X_{t_{n}})>2^{n}.$ Define a random variable $X$ by
$X=\sum_{n=1}^{\infty}2^{-n}X_{t_{n}}.$
Note that $X$ is in $L^{1}$ since $X\geqslant 0$ and $\mathbb{E}[X]=\sum_{n=1}^{\infty}\mathbb{E}[X_{t_{n}}]\leqslant 1.$ Finally, we note that the set of random variables $(X_{t_{n}})_{n\in\mathbb{N}}$ is comonotonic. Since any distortion risk measure is comonotonic-additive and positively homogeneous, we have
$\rho(X)=\sum_{n=1}^{\infty}\rho(2^{-n}X_{t_{n}})=\sum_{n=1}^{\infty}2^{-n}\rho(X_{t_{n}})\geqslant\sum_{n=1}^{\infty}1=\infty.$
This shows that $\rho$ is not finite on $L^{1}$.
(iii)$\Rightarrow$(i). Let $c=\lim_{t\downarrow 0}h(t)/t<\infty$. Since $h(t)/t$ is decreasing in $t$, we have $h(t)\leqslant ct$ for $t\in[0,1]$. Hence, $\rho_{h}(X)\leqslant c\mathbb{E}[X]$ for $X\in L^{1}_{+}$. ∎
Theorem 4 generalizes a special case of Proposition 1 of Wang et al. (2020), which states that if $h$ is concave and $\rho_{h}$ is finite on $L^{1}$, then the right derivative $h^{\prime}$ of $h$ has finite $L^{q}$-norm on $[0,1]$ with respect to the Lebesgue measure for all $q>0$. Theorem 4 further gives that $h^{\prime}$ has finite $L^{\infty}$-norm, which is a stronger condition.
Remark 3.
Theorem 4 (i)$\Leftrightarrow$(iii) can be derived from some results in the context of rearrangement-invariant Banach norms. In particular, we find Theorem 29 and Corollary 30 of Fröhlich and Williamson (2024) the most relevant. Translating their results on Banach norms into our setting of distortion risk measures, it is shown that for $h\in\mathcal{D}$ with $h(0+)=0$, if $h^{\prime}(0)<\infty$ then the induced norm by $\rho_{h}$ via $X\mapsto\rho_{h}(|X|)$ is equivalent to the $L^{1}$ norm; if $h^{\prime}(0)=\infty$ then the induced norm by $\rho_{h}$ is not equivalent to the $L^{1}$ norm. Our proof uses an explicit construction to show the implication (ii)$\Rightarrow$(iii), not covered by Fröhlich and Williamson (2024), and includes the case $h(0+)>0$. A similar observation can be made for Proposition 3 below, which characterizes the finiteness of law-invariant coherent risk measures on $L^{1}$.
Using Theorem 4, we can show that there does not exist a single $\rho_{h}$ that works for all sets $\mathcal{S}$ in Theorem 3.
Proposition 2.
There is no coherent distortion risk measure $\rho$ such that for all sets $\mathcal{S}\subseteq\mathcal{X}$,
$\sup_{X\in\mathcal{S}}\rho(|X|)<\infty~{}~{}\iff~{}~{}\mbox{$\mathcal{S}$ is uniformly integrable}.$
Proof.
First, by Theorem 4, if $\rho$ is finite on $L^{1}$, then it is expectation-dominated. In this case $\sup_{X\in\mathcal{S}}\rho(|X|)<\infty$ is merely $\sup_{X\in\mathcal{S}}\mathbb{E}[|X|]<\infty$, which is not sufficient for uniform integrability of $\mathcal{S}$. Next, by Theorem 4 again, if $\rho$ is not finite on $L^{1}$, then there exists $X\in L^{1}$ such that $\rho(|X|)=\infty$, but $\{X\}$ is uniformly integrable. Hence, uniform integrability of $\mathcal{S}$ does not guarantee $\sup_{X\in\mathcal{S}}\rho_{h}(X)<\infty.$
Example 6 (Integrated ES).
Define the integrated ES (IES) for a random variable $X$ as
$\mathrm{IES}(X)=\int_{0}^{1}\mathrm{ES}_{p}(X)\mathrm{d}p=\int_{0}^{1}\int_{0}^{q}\frac{1}{1-p}\mathrm{VaR}_{q}(X)\mathrm{d}p\mathrm{d}q=\int_{0}^{1}-\log(1-q)\mathrm{VaR}_{q}(X)\mathrm{d}q.$ (7)
The distortion function $h$ of $\mathrm{IES}$ is in $\mathcal{D}_{c}$ because $\lim_{t\downarrow 0}h(t)/t=-\lim_{t\downarrow 0}\log(t)=\infty.$ By Theorem 3, any set $\mathcal{S}$ of random variables on which $\mathrm{IES}$ is bounded is uniformly integrable. Proposition 1 of Wang et al. (2020) implies that $\mathrm{IES}$ is finite on $L^{1+\varepsilon}$ for any $\varepsilon>0$, and by Theorem 4, $\mathrm{IES}$ is not finite on $L^{1}$. One specific example of $X\in L^{1}$ with $\mathrm{IES}(X)=\infty$ is given below. Let $U$ be uniformly distributed on $[0,1/2]$ and let $X=U^{-1}(\log U)^{-2}$. We can compute
$\mathbb{E}[X]$ $=2\int_{0}^{1/2}(-\log u)^{-2}u^{-1}\mathrm{d}u=2\int_{\log 2}^{\infty}t^{-2}\mathrm{d}t=2(\log 2)^{-1}<\infty.$
By (7) and $\mathrm{VaR}_{1-q}(X)=2q^{-1}(\log q-\log 2)^{-2}$,
$\mathrm{IES}(X)=\int_{0}^{1}-\log u\frac{2}{u(\log u-\log 2)^{2}}\mathrm{d}u=\int_{0}^{\infty}\frac{2x}{(x+\log 2)^{2}}\mathrm{d}x=\infty.$
We can see that the set $\{X\}$ is uniformly integrable and $\mathrm{IES}(X)=\infty$.
4.3 Uniform integrability and coherent risk measures
Next, we turn to the more general class of law-invariant coherent risk measures. Let $\mathcal{R}$ be the set of all lower semicontinuous and law-invariant coherent risk measures on $L^{1}$ that are not finite on $L^{1}$. For instance, IES in Example 6 is in $\mathcal{R}$.
Theorem 5.
For any set $\mathcal{S}$ of random variables, the following are equivalent.
  1. (i)
    $\mathcal{S}$ is uniformly integrable;
  2. (ii)
    there exists $\rho\in\mathcal{R}$ such that $\rho(|X|)$ is bounded for $X\in\mathcal{S}$;
  3. (iii)
    there exists $\rho\in\mathcal{R}$ such that $\rho(X)$ and $\rho(-X)$ are bounded for $X\in\mathcal{S}$.
Proof.
Any coherent distortion risk measure $\rho_{h}$ is lower semicontinuous. By Theorem 4, $\rho_{h}\in\mathcal{R}$ if and only if $h\in\mathcal{D}_{c}$. Therefore, by using Theorem 3 and taking $\rho=\rho_{h}$, we get (i)$\Rightarrow$(ii)$\Rightarrow$(iii).
Next, we prove (iii)$\Rightarrow$(i). We first show that for any coherent distortion risk measure $\rho_{h}$ and $c>0$, we have
[…]
A weaker version of this equivalence was used in the proof of Theorem 4. Suppose $\lim_{t\downarrow 0}h(t)/t\leqslant c$. Since $h(t)/t$ is decreasing in $t$, we know $h(t)\leqslant ct$ for $t\in[0,1]$. Hence, $\rho_{h}(X)\leqslant c\mathbb{E}[X]$ for $X\in L^{1}_{+}$. Conversely, if $\rho_{h}\leqslant c\mathbb{E}$ on $L^{1}_{+}$, then $\rho_{h}(B_{t})/t\leqslant c$, where $B_{t}$ is a Bernoulli random variable with $\mathbb{E}[B_{t}]=t$. This implies $h(t)/t$ is bounded above by $c$. Since $h(t)/t$ is decreasing in $t$, it has a finite limit bounded by $c$ as $t\downarrow 0$.
Note that a lower semicontinuous and law-invariant coherent risk measure $\rho$ on $L^{1}$ admits a Kusuoka representation (Kusuoka (2001); see Filipović and Svindland (2012) for its extension to $L^{1}$), that is,
$\rho=\sup_{h\in\mathcal{H}_{\rho}}\rho_{h}\mbox{~{} for a set $\mathcal{H}_{\rho}$ of concave distortion functions}.$ (9)
Since $\rho$ is not finite on $L^{1}$, it cannot be expectation-dominated; this follows by the same argument of (i)$\Rightarrow$(ii) in the proof of Theorem 4. That is, for every $c>0$, $\rho\leqslant c\mathbb{E}$ on $L^{1}_{+}$ does not hold. Hence, for each $n$, there exists $h_{n}\in\mathcal{H}_{\rho}$ such that $\rho_{h_{n}}\leqslant 2^{n}\mathbb{E}$ on $L^{1}_{+}$ does not hold, and by using (8), this implies $\lim_{t\downarrow 0}h_{n}(t)/t>2^{n}.$ Write $g=\sum_{n=1}^{\infty}2^{-n}h_{n}$. Since $g$ is a convex combination of concave distortion functions, it is also a concave distortion function. Moreover, we can compute (via Fubini’s theorem)
$\lim_{t\downarrow 0}\frac{g(t)}{t}=\sum_{n=1}^{\infty}2^{-n}\lim_{t\downarrow 0}\frac{h_{n}(t)}{t}\geqslant\sum_{n=1}^{\infty}1=\infty.$
Therefore, $g\in\mathcal{D}_{c}$. Moreover, $\rho\geqslant\rho_{g}$ since $\rho\geqslant\rho_{h_{n}}$ for each $n\in\mathbb{N}$. Hence, the boundedness of $\rho(X)$ and $\rho(-x)$ for $X\in\mathcal{S}$ implies that $\rho_{g}(X)$ and $\rho_{g}(-X)$ are both bounded above; they are also bounded below because $\rho_{g}(X)+\rho_{g}(-X)\geqslant 0$ for all $X\in L^{1}$. By using Theorem 3, we know that $\mathcal{S}$ is uniformly integrable. ∎
Let $\rho$ be a lower semicontinuous and law-invariant coherent risk measure on $L^{1}$. In the direction (iii)$\Rightarrow$(i) of Theorem 5, it is required that $\rho$ is not finite on $L^{1}$. It may be natural to ask whether this requirement is also necessary. Clearly, if $\rho$ is expectation-dominated, then $\sup_{X\in\mathcal{S}}\rho(|X|)<\infty$ for any $\mathcal{S}\subseteq L^{1}_{+}$ with bounded expectation, which is not sufficient for uniform integrability. Therefore, for the implication (iii)$\Rightarrow$(i), the choice of $\rho$ is necessarily not expectation-dominated. It turns out that this is equivalent to requiring $\rho$ to be not finite. The next result on the finiteness of such risk measures extends Theorem 4 to the larger class of law-invariant coherent risk measures. For other results on spaces that are finite for law-invariant risk measures, see Liebrich and Svindland (2017).
Proposition 3.
A law-invariant coherent risk measures on $L^{1}$ is finite if and only if it is expectation-dominated.
Proof.
The “if" statement follows from the same argument of (i)$\Rightarrow$(ii) in the proof of Theorem 4. Below we show the “only if" statement. Note that any law-invariant coherent risk measure that is finite on $L^{1}$ admits a Kusuoka representation in the form of (9). From the arguments in the proof of Theorem 5, we see that $\rho\geqslant\rho_{g}$ for some $g\in\mathcal{D}_{c}$. By Theorem 4, this implies that $\rho_{g}$ is not finite on $L^{1}$. Hence, we conclude that $\rho$ is also not finite on $L^{1}$. ∎
Let $\rho$ be a lower semicontinuous and law-invariant coherent risk measure. Proposition 3 implies that the set $\mathcal{R}$ contains precisely such $\rho$ that are not expectation-dominated. Moreover, the non-finiteness of $\rho$ on $L^{1}$ is necessary and sufficient for the implication (iii)$\Rightarrow$(i).
5 Two consequences of the main results
In this section we present two applications, both following directly from the condition of uniform integrability. The first corollary concerns a weak law of large numbers.
Corollary 1.
Let $(X_{n})_{n\in\mathbb{N}}$ be a sequence of pairwise independent random variables with zero mean and $Y_{n}=(X_{1}+\dots+X_{n})/n$ for each $n\in\mathbb{N}$. If $\sup_{n\in\mathbb{N}}(\rho(X_{n})\vee\rho(-X_{n}))<\infty$ for some $\rho\in\mathcal{R}$, then $Y_{n}\to 0$ in probability as $n\to\infty$.
Proof.
For a sequence of pairwise independent random variables with zero mean, uniform integrability ensures the weak law of large numbers; see Landers and Rogge (1987). The statement then follows from Theorem 5. ∎
Remark 4.
It is known that, for a sequence of independent random variables with zero mean, uniform integrablity is not sufficient for the strong law of large numbers; see the example in Landers and Rogge (1987). Since our condition in Corollary 1 is equivalent to uniform integrability, we cannot change the statement of convergence in probability to that of almost sure convergence. A remaining question is to find a stronger condition on $\rho$ than $\rho\in\mathcal{R}$ such that
[…]
for an independent (or pairwise independent) sequence $(X_{n})_{n\in\mathbb{N}}$ with zero mean.
The second corollary concerns the convergence of ES under convergence in distribution.
Corollary 2.
Suppose that a sequence $(X_{n})_{n\in\mathbb{N}}$ converges to $X\in L^{1}$ in distribution and satisfies $\sup_{n\in\mathbb{N}}(\rho(X_{n})\vee\rho(-X_{n}))<\infty$ for some $\rho\in\mathcal{R}$. Then $\mathrm{ES}_{p}(X_{n})\to\mathrm{ES}_{p}(X)$ for all $p\in(0,1)$.
Proof.
Let $U$ be a uniform random variable on $[0,1]$. For $n\in\mathbb{N}$, denote by $Q_{n}(t)=\mathrm{VaR}_{t}(X_{n})$ and $Q(t)=\mathrm{VaR}_{t}(X)$ for $t\in(0,1)$. Note that $Q_{n}(U)\buildrel\mathrm{d}\over{=}X_{n}$ and $Q(U)\buildrel\mathrm{d}\over{=}X$. Using Theorem 5, we know that $\{X_{n}:n\in\mathbb{N}\}$ is uniformly integrable, so $\{Q_{n}(U):n\in\mathbb{N}\}$ is also uniformly integrable. The convergence $X_{n}\to X$ in distribution as $n\to\infty$ implies that $Q_{n}(U)\to Q(U)$ in probability. This and uniform integrability of $\{Q_{n}(U):n\in\mathbb{N}\}$ guarantee that $Q_{n}(U)\to Q(U)$ in $L^{1}$. Hence, $\int_{0}^{1}|Q_{n}(u)-Q(u)|\mathrm{d}u\to 0,$ which implies $\int_{p}^{1}|Q_{n}(u)-Q(u)|\mathrm{d}u\to 0.$ Therefore, by law invariance of ES, we have $\mathrm{ES}_{p}(X_{n})=\mathrm{ES}_{p}(Q_{n}(U))\to\mathrm{ES}_{p}(Q(U))=\mathrm{ES}_{p}(X).$
6 Conclusion
The main contribution of this paper is to establish a connection between boundedness of risk measure values and uniform integrability. As a convenient technical tool, we obtained an upper bound on the folding score of each distortion risk measure (Theorem 1). Three different sets of equivalent conditions for uniform integrability are obtained, via ES (Theorem 2), via distortion risk measures (Theorem 3) and via law-invariant coherent risk measures (Theorem 5). Conditions in these results are stated both for the absolute value of the random variables involved and for the random variables themselves, facilitating easy interpretation in risk management. These results form a bridge between two important concepts, one in probability theory and one in financial mathematics. Moreover, they highlight the symmetric roles played by the mappings $X\mapsto\rho_{h}(X)$ and $X\mapsto\mathbb{E}[\phi(X)]$, well known in decision theory (Yaari (1987)). A mathematical duality between the distributional transforms underlying these two classes of mappings, in the sense that one class is characterized by commutation with the other, is recently obtained by Chambers et al. (2023).
Some remaining questions concern boundedness of the folding score of general law-invariant coherent risk measures, and conditions on the “testing" risk measure for a strong law of large numbers. These two questions are discussed in Remarks 2 and 4.
Acknowledgements
The authors thank Felix Liebrich for very helpful discussions and comments. RW acknowledges financial support from the Natural Sciences and Engineering Research Council of Canada (RGPIN-2018-03823 and CRC-2022-00141).